OpenStudy (anonymous):

Chico walks 3/5 mile to the library. Nikki walks 1/4 the distance that Chico walks to the library. How far does Nikki walk?

OpenStudy (mathmale):

How far does Chico walk? that's given; it's 3/5 mile, or \[\frac{ 3mile }{ 5 }.\]

OpenStudy (mathmale):

We compare the distances that these two peoople walk. Nikki walks only a quarter as far as does Chico. This calls for us to multiply\[\frac{ 3mile }{ 5 }.by.\frac{ 1 }{ 4 }\]

OpenStudy (mathmale):

Try it, please. show your work.

OpenStudy (anonymous):

3/20

OpenStudy (anonymous):

o is the answer is 36

OpenStudy (mathmale):

the answer couldn't possibly be 36. Chico walks only three-fifths of a mile (0.6 mile). Try again.

OpenStudy (anonymous):

im pretty sure its 3/20ths a mile

OpenStudy (anonymous):

o ok